OpenStudy (anonymous):

Write the equation of a line with a slope of -3 that passes through (-2,5)

OpenStudy (gorv):

standard eqution of line y=mx+c m=slope put the point in equation and find the value of c

OpenStudy (gorv):

will u give it a try

OpenStudy (08surya):

y-5=-3(x+2)

OpenStudy (08surya):

y-y1=m(x-x1) put m=slope and x1,y1 is the point at which line is passing